# Modern approaches to building effective interpretable models of the property market using machine learning

### v1

@@ 4 Formation and selection of explanatory variables
According to methodological rules of mass valuation of real estate, we can use for explanatory variables such information as general characteristics of an object (object’s purpose, area, object’s name, wall material, year build, extension of the land parcel, kind of soil and so on); information about a location of an object (a distance to the roads, to the water objects, to a recreation area, a city center and so on); information about engineering infrastructure (a class of linear object, a distance to a power line, a rated power of a linear object and so on); other matter (a price level of a commodity bundle, commodity circulation per 1 person, a distance to extractive industry territory) (Ministry for Economic Development of Russia, 2018).
@@
General information about the object is contained in the Unified State Register of Real Estate and in accordance with the law is given by the ministry to the certified organization which makes evaluation. Socioeconomic data can be found in the annual reviews of the development of the territory. However such information can be useful only for large territories that contain a few municipal districts or a few towns. Geographical features can be extracted with the help of mapping software – NextGIS and Arcgis. Also, we can get such information using programming tools, for example, Python and open non-profit mapping source – OpenStreetMap. Using the library “osmnx” from Python we can get all kinds of socially significant facilities – nursery schools, schools, hospitals, shops, malls, and so on. Then we can calculate the straight-line distances to these objects or, for example, count the number of such objects in some radius. We can use this data as explanatory variables for a predictive model. Also libraries “osmnx” and “networkx” allow loading road network graphs. So we have the opportunity to calculate the distance by the roads. This calculation improves the accuracy of the variables. It can be important when we have a territory with irregular coastlines (Figure 6) or deal with islands. In the last case, we can use adjusting factors reflecting the accessibility of the island.
@@
Also, we can aggregate some variables into one (as explained below in the discussion of the correlation matrix) or use conceptions from the related fields, for example, graph theory. Thus we created the factor “Development of the road network” that helped to estimate the location of object in relation to transport accessibility. If the object has many crossroads and roads nearby then this object is more attractive and therefore more expensive. This idea is more important for big cities. For the derivation of the factor “Development of the road network”, we used the road network graphs and the conception of centrality. The factor of centrality in graph theory and road network analysis allows to find the most important nodes of a graph (Newman, 2010). Important nodes are points that have the highest number of shortest routes. For the calculation of centrality, we consider the points of crossroads and the endpoints of roads as nodes of a graph and roads between these points as graph edges. In these terms the importance of a node means how many shortest roads between two points go through this node. When we calculate the shortest routes we naturally associate the road distance between two points with weights of the corresponding graph edges. We use the Python library “networkx” to calculate the centrality. This library allows to calculate the centrality in different ways. In this paper, we use harmonic centrality (Marchiori and Latora, 2000) because it is simple for calculations and understanding. The harmonic centrality of a node is the sum of inverse lengths of the shortest routes for each node to a particular node. It has a perfect intuitive interpretation – smaller distances give a higher contribution to centrality. Mathematically the harmonic centrality can be written as:
@@
where - the node for which the centrality is calculated; - the other nodes of the graph; - the shortest route between and . The shortest route between two points is the route for which the sum of weights of constituting edges is minimal.
@@
The roads can have bi-directional and unidirectional traffic. Because the graph of the road network is an oriented graph, it means that every edge of the graph has a direction. We should take this fact into account when calculating the centrality. We calculated the harmonic centrality for the road network of Vladivostok (Figure 7 (a)).
@@
second step: consider this result as an estimation of the centrality of the object. -
@@
where - estimation at the point ; - distance from point to point ; - basis function that depends on the distance; - weight coefficients. Coefficient defines algebraic sign of an element and its contribution.
@@
The weight coefficients are derived from the condition that at known points the results of interpolation must coincide with known values, i.e. the interpolation model in all points must estimate values equal to . Thus we need to solve the system of linear equations with unknown variables.
To do this we use the function RBFInterpolator from Python library ”SciPy”. The results of interpolation for Vladivostok are shown in Figure 7 (b). Here for the centrality of the objects we use the term “Development of the road network” underlining our mathematical interpretation of this concept.
If we want to build a model for an inhomogeneous area (for example if territory contains a few towns or villages) then we can try to introduce new variables, using aggregation of existing variables. For example, we created the factor “Population of a residential place divided on the distance to Vladivostok by the roads”. On the one hand, the increasing of distance between the residential place and Vladivostok’s administration negatively affects the per square meter price; on the other hand, the population of the residential place positively affects the per square meter price. These considerations led us to create the factor that aggregates two variables – the population of a residential place and the distance to Vladivostok by the roads. Econometrically this factor should have a positive correlation with the per square meter price – with an increase in the population of residential place and a decrease inf the distance to Vladivostok the per square meter price increases. Figure 8(a) shows the relation between this factor and the average value of per square meter price over a building. We can notice that data contains positive a linear correlation however we also can notice that there are quite a lot of points when the aggregate factor has low values. Because of this, we tried to raise this factor to different powers. Figure 8(b) shows the relation between this factor raised to 1/3 power and the average value of per square meter price over a building. Such transformation gives a better value of correlation.
@@
When we study the matrices of correlation (Figure 9, Figure 10) we are interested not only in the relationship between explanatory variables and dependent variable but also in the relationships between explanatory variables with each other. The first model that we tried was the linear regression, because of its interpretability. The linear model has a set of assumptions, one of them is the absence of multicollinearity in data. Multicollinearity is a situation when two or more variables have a significant correlation with one another (). In this case, the inclusion of them in the model becomes unpractical. Because of this multicollinearity should be removed from data before modeling. We can achieve it in several ways: we can keep only factors that have a more significant correlation with the dependent variable; we can drop factors with the wrong sign of correlation with the dependent variable, which means that such factors have incorrect economic substance; finally, we can integrate several factors into one employing the principal component analysis (PCA) (Ajvazyan, 2010). This method is useful when several factors have similar economic substance. For example, the distance to school, nursery school, and supermarkets can be unified into one factor which means the distance to socially significant facilities (alternatively instead of distance we can use the number of objects). The reason for such aggregation is that these factors have significant relation with one another, but if we drop one of them we will lose some information. Instead, we incorporate them into the principal component which includes the dispersion of all these factors allowing to save more information about our variables. It is worth noting that the majority of considered features reflect spatial characteristics of data. The inclusion of such features in the model is critically important for building an effective model of spatial data (Jahanshiri et al., 2011) like land parcels and flats.
@@
Previously we noted that before including categorical factors into a model we plot probability density functions of a per square meter price. A similar method of analysis would be drawing of a box-plot diagram. Traditionally the categorical factors are included in a model as binary features and one should pay attention not only to average values of per square meter price for each group but also to the representativity of the group in the data. For example, Figure 11 shows the probability density function for the factor “Wall material” and we see that flats in timber buildings have a per square meter price far less than the other materials. For this reason, we can’t include this binary feature in a model due to its nonrepresentativity. A possible solution to this problem can be the introduction of a correction factor reducing the per square meter price of such objects after getting of estimation of a model.
@@
After completing all preparative steps – removing multicollinearity; dropping the factors without the economic substance; and dropping the non-significant features, we can move to the question of selection of price-forming factors that we may include in the model. There are different ways to solve this problem (Ajvazyan, 2010) – we can try all possible combinations of factors and choose the best model taking into account some requirements (minimizing errors or maximizing ); we can choose a model that meets the initial conditions for its using (for example for ordinary least squares such conditions are unbiasedness, consistency and effectiveness of estimations); also we can use algorithms for automatic feature selection – Forward selection or Backward elimination (Narisetty, 2020).
@@
For segment “Land parcels” where the correlation of features with dependent variable has sufficiently high values (Figure 9) we applied Forward selection of features and also we took into account the value of and realization of the basic conditions for ordinary least squares method (Ajvazyan, 2010; Kutner, 2005) (Durbin–Watson statistic, Jarque-Bera test, F-test).
Number of factors | D–W stat | J-B (prob) | F-stat (prob) | |
1 | 0,636865 | 1,238017 | 0,000005 | 1,173483e-207 |
2 | 0,712070 | 1,371189 | 0,072956 | 3,926066e-253 |
3 | 0,749263 | 1,564183 | 0,175133 | 1,023501e-279 |
4 | 0,752834 | 1,564116 | 0,073893 | 2,622868e-281 |
5 | 0,763181 | 1,609770 | 0,002961 | 1,090438e-288 |
6 | 0,773200 | 1,566779 | 0,004079 | 3,415025e-296 |
@@
From Table 2 one can see that in this particluar case, 4 factors are the optimal choice for modeling linear regression taking into account the conditions for using ordinary least squares.
@@ 5 Modeling
In this paper, we show the results of the prediction of the cadastral value of objects located in Primorsky Krai – a sufficiently large region of the Russian Federation. The area of this region is 165 thousand square kilometers. The region has a capital city – Vladivostok (with a population of about 600 thousand people), towns (with a population of not more than 150 thousand people), and villages. All these territories have different property markets, price levels, and quality of life, that depend on economic and territorial characteristics. The methodological rules (Ministry for Economic Development of Russia, 2018) recommend splitting the territory on the base of administrative boundaries of regions taking into account the population. This recommendation can be applied to objects that are located on homogeneous territory, for example for flats that are located in one city. However, this recommendation works poorly for objects that are spread over a large area e.g. land parcels. They occupy a large heterogeneous area – the city and the suburbs and often form allotment societies (also called gardeners’ non-commercial partnerships). Gardeners’ non-commercial partnerships are a collective form of gardening. The members perform their activities on an individual basis. Activities of such partnerships are regulated by the federal law. The partnerships can have their regulations, convoke meetings, and appoint a chairman. Similar organizations exist in many countries, for example, Germany, Sweden, Poland, and the USA. In every country, there are own procedures for the registration and obligation to pay a tax. In Russia, such partnerships must pay taxes, in particular land tax (Seregina, 2019) – the tax paid by physical persons or organizations that have property right of land plot or right of land plot perpetual usage or right of lifetime inheritable possession. Tax amount is defined from cadastral value and taxation rate. Such an allotment society can include different administrative regions. In such cases, it is better to use a grouping of objects employing clusterization algorithms instead of using formal administrative borders. Most of the algorithms are based on spatial relations between objects. Because of this reason, for the segment “Land parcels” we used a k-means clustering algorithm (MacQuenn, 1967), that takes into account the spatial positions of objects. This method is simple for understanding and for realization. It recognizes the accumulation of points and creates the clusters in such a way that the distance to the center of the cluster for every point must be minimal. By using this algorithm we are more confident that neighboring points belong to one cluster.
All further calculations of modeling were done for the capital region – Vladivostok and its suburbs.
@@
First of all, we plotted the probability density function of the dependent variable (per square meter price) and noticed that it is not normal (Figure 12(a)). Although statistic and machine learning theories don’t have the strict rule requiring the normality of the probability density function of the dependent variable, usually analysts prefer to work with data distributed approximately normally. Because of this reason, we used a natural logarithm of the per square meter price (Figure 12(b)). Also such transformation of data guarantees that prediction of modeling will have only positive values.
@@
We start the modeling by trying the simple and interpretable model – the linear regression. As we mentioned above for the selection of features we apply Forward selection using a correlation matrix (Figure 9). The data was split into two parts – a training set (70% of data) and a testing set (30% of data). Figure 13 shows the result of the training of the model.
Using Figure 13 we can check the model on basic requirements for linear regression. We see that the coefficients of variables are significant (); in general, the equation of regression is significant (); the coefficients of variables have the right sign according to economic substance; model residuals have normal distribution () and don’t have autocorrelation (); adjusted coefficient of determination is significant () and it is not differ sufficiently from .
Also factors included in the model don’t have multicollinearity because the value of the variance inflation factor (VIF) is less than 10 for each factor (Shrestha, 2020) (Table 3).
Factors | |
Distance to the water objects (straight-line distance) | 1,463 |
Distance to the seacoast (straight-line distance) | 1,320 |
Distance to the town hall (straight-line distance) | 1,936 |
@@
Aggregate: Distance to the hospitals (straight-line distance) & Distance to the closest nursery school (straight-line distance) & Distance to the closest school (straight-line distance) | 1,424 |
@@
Figure 14 shows the distribution of model residuals. This distribution is approximately normal which satisfies the assumptions of linear regression. Also, we tested the normality of residuals using the library “SciPy”. This test compares calculated with the significance threshold (for this paper we chose ). If calculated is more than we can assume that the distribution of model residuals is normal (in our case ).
@@
Also, we created the scatter plot values of residuals from their numerical order. From Figure 15 we can see that the data does not contain any dependencies. It means that residuals are homoscedastic and the other assumption of linear regression is satisfied.
@@
On the base of the described results, we can state that our model is effective and interpretable. As we’ve already said the key factor for the identification of the price of a land parcel is its location. We see that linear regression does relatively well but the price of an object can include latent (local-territorial) characteristics. In order to take this into account we tried to apply a powerful instrument dealing with spatially distributed data – geostatistics. Geostatistical methods allow to model the distribution of objects in space. The central task of geostatistics is to reconstruct a studied phenomenon on the base of the values calculated in a limited number of points (Dem’yanov and Savel’eva, 2010). Mathematically we can contemplate this problem as an interpolation problem. For getting the best spatial estimations kriging methods are used. Estimations obtained by kriging have a minimal variation of error. Thus kriging is the best estimator in a class of linear interpolators (Dem’yanov and Savel’eva, 2010). There are two basic conditions for using these methods – data should have spatial continuity and stationarity (Kovalevskiy, 2012). The first condition means that two points located close together should have rather similar values than points separated more. The second condition means that in some locality the studied phenomenon must be spatially homogeneous.
@@
We can check the condition of stationarity using the probability density function of a variable, in our case we analyze the distribution of residuals of our linear model. If the histogram is similar to the Gaussian distribution then in all likelihood there is stationarity (statistical homogeneity). Strictly speaking, this criteria is not a sufficient condition for confirmation of stationarity but it is strong enough and sufficient condition for practical work (Kovalevskiy, 2012). Figure 14 and the result of the Jarque–Bera test confirm the stationarity in linear model residuals. It’s worth noting that before checking data on stationarity we need to examine whether data contains a trend. In our case, we removed the trend from data using linear regression and then we modeled its residuals with geostatistical methods. The trend in data can distort the variogram – an instrument for estimation of spatial correlations in data (Dem’yanov and Savel’eva, 2010; Isaaks, 1989). The variogram model is a function describing a dependence between the studied spatial variable and the distance. If data contains a trend then a variogram would reproduce a large-scale trend that leads to missing a correlation of the observed variable on a small scale.
@@
The second condition for implementing kriging – the spatial continuity. We checked it by plotting the correlation of points as a function of the distance between them. From Figure 17 we see that with increasing distance between points, the correlation decreases. Therefore this criteria is satisfied.
Thus the final value of the per square meter price is the sum of the trend that we got using the linear regression and the prediction of linear model residuals that we got using kriging.

### v2

@@ 4 Formation and selection of explanatory variables
According to methodological rules of mass valuation of real estate, we can use as explanatory variables such information as general characteristics of an object (object’s purpose, area, object’s name, wall material, year built, extension of the land parcel, kind of soil and so on); information about a location of an object (a distance to the roads, to the water objects, to a recreation area, a city center and so on); information about engineering infrastructure (a class of linear object, a distance to a power line, a rated power of a linear object and so on); other matter (a price level of a commodity bundle, commodity circulation per 1 person, a distance to extractive industry territory) (Ministry for Economic Development of Russia, 2018).
@@
General information about the object is contained in the Unified State Register of Real Estate and in accordance with the law is given by the ministry to the certified organization which makes evaluation. Socioeconomic data can be found in the annual reviews of the development of the territory. However such information can be useful only for large territories that contain a few municipal districts or a few towns. Geographical features can be extracted with the help of mapping software – NextGIS and ArcGIS. Also, we can get such information using programming tools, for example, Python and open non-profit mapping source – OpenStreetMap (OpenStreetMap contributors, 2017). Using the library “OSMnx” (Boeing, 2025) from Python we can get all kinds of socially significant facilities – nursery schools, schools, hospitals, shops, malls, and so on. Then we can calculate the straight-line distances to these objects or, for example, count the number of such objects in some radius. We can use this data as explanatory variables for a predictive model. Also libraries “OSMnx” and “NetworkX” (Hagberg et al., 2008) allow the loading of road network graphs. So we have the opportunity to calculate the distance by the roads. This calculation improves the accuracy of the variables. It can be important when we have a territory with irregular coastlines (Figure 6), or mountain ranges, or deal with islands. In the last case, we can use adjusting factors reflecting the accessibility of the island.
@@
Also, we can aggregate some variables into one (as explained below in the discussion of the correlation matrix) or use conceptions from the related fields, for example, graph theory. Thus we created the factor “Development of the road network” that helped to estimate the location of an object in relation to transport accessibility. If the object has many crossroads and roads nearby then this object is more attractive and therefore more expensive. This idea is more important for big cities. For the derivation of the factor “Development of the road network”, we used the road network graphs and the conception of centrality. The factor of centrality in graph theory and road network analysis allows to find the most important nodes of a graph (Newman, 2010). Important nodes are points that have the highest number of shortest routes. For the calculation of centrality, we consider the points of crossroads and the endpoints of roads as nodes of a graph and roads between these points as graph edges. In these terms the importance of a node means how many shortest roads between two points go through this node. When we calculate the shortest routes we naturally associate the road distance between two points with weights of the corresponding graph edges. We use the Python library “networkx” to calculate the centrality. This library allows to calculate the centrality in different ways. In this paper, we use harmonic centrality (Marchiori and Latora, 2000) because it is simple for calculations and understanding. The harmonic centrality of a node is the sum of inverse lengths of the shortest routes from each node to this node. It has a perfect intuitive interpretation – smaller distances give a higher contribution to centrality. Mathematically the harmonic centrality can be written as:
@@
where is the node for which the centrality is calculated; is the other nodes of the graph; is the length of the shortest route between and . The shortest route between two points is the route for which the sum of weights of constituting edges is minimal.
@@
The roads can have bi-directional and unidirectional traffic. Because the graph of the road network is an oriented graph, it means that every edge of the graph has a direction so the inversion of end points of the route potentially could change its length. We should take this fact into account when calculating the centrality. We calculated the harmonic centrality for the road network of Vladivostok (Figure 7 (a)).
@@
second step: consider its centrality as an estimation of the centrality of the object. -
@@
where is the estimation of centrality at the point ; is the distance from point to point ; is the basis function that depends on the distance; are weight coefficients. Coefficient defines algebraic sign of an element and its contribution.
@@
The weight coefficients are derived from the condition that at known points the results of interpolation must coincide with known values, i.e. the interpolation model in all points must provide values equal to . Thus we need to solve the system of linear equations with unknown variables.
To do this we use the function RBFInterpolator from Python library “SciPy” (Virtanen et al., 2020; SciPy team, ). The results of interpolation for Vladivostok are shown in Figure 7 (b). Here for the centrality of the objects we use the term “Development of the road network” underlining our mathematical interpretation of this concept.
@@
If we want to build a model for an inhomogeneous area (for example if territory contains a few towns or villages) then we can try to introduce new variables, using aggregation of existing variables. For example, from Figure 8 (a) one can see that independent variables “: Population of a human settlement” and “: Distance to Vladivostok (by roads)” have quite significant correlation with the dependent variable “: Per square meter price” around 0.5. But we can achieve even higher correlation with by creating the new factor “Population of a residential place divided by the distance to Vladivostok by roads”. On the one hand, the increasing of distance between the residential place and Vladivostok’s administration negatively affects the per square meter price; on the other hand, the population of the residential place positively affects the per square meter price. These considerations led us to create the factor that aggregates two variables – the population of a residential place and the distance to Vladivostok by roads. Econometrically this factor should have a positive correlation with the per square meter price – with an increase in the population of residential place and a decrease in the distance to Vladivostok the per square meter price increases. Figure 8 (b) shows the relation between this factor and the value of per square meter price of flat.
@@
When we study the matrices of correlation (Figure 9, Figure 10) we are interested not only in the relationship between the explanatory variables and the dependent variable but also in the relationships between explanatory variables. The first model that we tried was the linear regression, because of its interpretability. The linear model has a set of assumptions, one of them is the absence of multicollinearity in data. Multicollinearity is a situation when two or more variables have a significant correlation with one another ( (Kozak, 2009; Gujarati and Porter, 2009)). In this case, the inclusion of them in the model becomes unpractical. Because of this multicollinearity should be removed from data before modeling. We can achieve it in several ways: we can keep only factors that have a more significant correlation with the dependent variable; we can drop factors with the wrong sign of correlation with the dependent variable, which means that such factors have incorrect economic substance; finally, we can integrate several factors into one employing the principal component analysis (PCA) (Ajvazyan, 2010). This method is useful when several factors have similar economic substance. For example, the distance to school, nursery school, and supermarkets can be unified into one factor which means the distance to socially significant facilities (alternatively instead of distance we can use the number of objects). The reason for such aggregation is that these factors have significant relation with one another, but if we drop one of them we will lose some information. Instead, we incorporate them into the principal component which includes the dispersion of all these factors allowing to save more information about our variables. It is worth noting that the majority of considered features reflect spatial characteristics of data. The inclusion of such features in the model is critically important for building an effective model of spatial data (Jahanshiri et al., 2011) like land parcels and flats.
@@
Previously we noted that before including categorical factors into a model we plot probability density functions of a per square meter price. A similar method of analysis would be drawing of a box-plot diagram (Tukey, 2000). Traditionally the categorical factors are included in a model as binary features and one should pay attention not only to average values of per square meter price for each group but also to the representativity of the group in the data. For example, Figure 11 shows the probability density function for the factor “Wall material” and we see that flats in timber buildings have a per square meter price far less than the other materials. For this reason, we can’t include this binary feature in a model due to its non-representativity. A possible solution to this problem can be the introduction of a correction factor reducing the per square meter price of such objects after getting of the estimation of a model.
@@
After completing all preparative steps – removing multicollinearity; dropping the factors without the economic substance; and dropping the non-significant features, we can move to the question of selection of price-forming factors that we may include in the model. There are different ways to solve this problem (Ajvazyan, 2010) – we can try all possible combinations of factors and choose the best model taking into account some requirements (minimizing errors or maximizing ); we can choose a model that meets the initial conditions for its usage (for example for ordinary least squares such conditions are unbiasedness, consistency and effectiveness of estimations (Gujarati and Porter, 2009)); also we can use algorithms for automatic feature selection – Forward selection or Backward elimination (Narisetty, 2020).
@@
For the segment “Land parcels” where the correlation of features with dependent variable has sufficiently high values (Figure 9) we applied Forward selection of features and also we took into account the value of (Gujarati and Porter, 2009) and realization of the basic conditions for ordinary least squares method (Ajvazyan, 2010; Kutner, 2005): residuals should not have the significant autocorrelation which corresponds to the Durbin–Watson statistic to be in range from 1.5 to 2.5 where the value 2.0 means that residuals are completely uncorrelated (Field, 2012); residuals should be normally distributed which means they should pass the Jarque-Bera test with -value(JB) 0.05 (Gujarati and Porter, 2009) that corresponds to do not rejecting the null hypothesis of normality; the equation of regression should be significant with -value(-statistic) 0.01 (Gujarati and Porter, 2009) meaning we should reject the null hypothesis that all coefficients of the model except the intercept are zero.
Number of factors | Durbin-Watson | -value(Jarque-Bera) | -value(-statistic) | |
1 | 0.634 | 1.238 | 0.000005 | 1.173483e-207 |
2 | 0.712 | 1.371 | 0.073 | 3.926066e-253 |
3 | 0.749 | 1.564 | 0.175 | 1.023501e-279 |
4 | 0.753 | 1.564 | 0.074 | 2.622868e-281 |
5 | 0.763 | 1.610 | 0.003 | 1.090438e-288 |
6 | 0.773 | 1.567 | 0.004 | 3.415025e-296 |
@@
From Table 2 one can see that in this particular case, 4 factors are the optimal choice for modeling linear regression taking into account the conditions for using ordinary least squares. The another option is to select 3 factors - it also satisfies the statistical tests, has the similar . But all other things being equal we think it is better to keep more factors.
@@ 5 Modeling
In this paper, we show the results of the prediction of the cadastral value of objects located in Primorsky Krai – a sufficiently large region of the Russian Federation. The area of this region is 165 thousand square kilometers. The region has a capital city – Vladivostok (with a population of about 600 thousand people), towns (with a population of not more than 150 thousand people), and villages. All these territories have different property markets, price levels, and quality of life, that depend on economic and territorial characteristics. The methodological rules (Ministry for Economic Development of Russia, 2018) recommend splitting the territory on the base of administrative boundaries of municipalities taking into account the population. This recommendation can be applied to objects that are located on homogeneous territory, for example for flats that are located in one city. However, this recommendation works poorly for objects that are spread over a large area e.g. land parcels. They occupy a large heterogeneous area – the city and the suburbs, and often form allotment societies (also called gardeners’ non-commercial partnerships). Gardeners’ non-commercial partnerships are a collective form of gardening. The members perform their activities on an individual basis. Activities of such partnerships are regulated by the federal law. The partnerships can have their regulations, convoke meetings, and appoint a chairman. Similar organizations exist in many countries, for example, Germany, Sweden, Poland, and the USA. In every country, there are its own procedures for registration and obligation to pay taxes. In Russia, such partnerships must pay taxes, in particular the land tax (Seregina, 2019) – the tax paid by individuals or legal entities that have the property right of land plot or the right of land plot perpetual usage or the right of lifetime inheritable possession. Tax amount is defined from the cadastral value and taxation rate. Such an allotment society can include different administrative regions. In such cases, it is better to use a grouping of objects employing clusterization algorithms instead of using formal administrative borders. Most of the algorithms are based on spatial relations between objects. Because of this reason, for the segment “Land parcels” we used a k-means clustering algorithm (MacQuenn, 1967), that takes into account spatial positions of objects. This method is simple for understanding and for realization. It recognizes the accumulation of points and creates the clusters in such a way that the distance to the center of the cluster for every point must be minimal. By using this algorithm we are more confident that neighboring points belong to one cluster. We used Scikit-learn realization of this algorithm.
All further calculations were done for the capital region – Vladivostok and its suburbs.
@@
First of all, we plotted the probability density function of the dependent variable (per square meter price) and noticed that it is not normal (Figure 12 (a)). Although there is no requirement on the normality of the dependent variable, analysts tend to prefer working with the data distributed approximately normally. Because of this reason, we used a natural logarithm of the per square meter price (Figure 12 (b)). Also such transformation of data guarantees that prediction of modeling will have only positive values.
OLS Regression Results | ||||
Dep. variable: | ||||
: | 0.761 | |||
: | 0.760 | |||
-statistic: | 740.8 | |||
-value (-statistic): | 1.27-287 | |||
coef | std err | t | ||
7.3210 | 0.012 | 627.655 | ||
0.1070 | 0.014 | 7.585 | ||
-0.1642 | 0.013 | -12.254 | ||
-0.3719 | 0.016 | -22.918 | ||
-0.1153 | 0.009 | -12.876 | ||
is Natural logarithm of per square meter price | ||||
is Distance to the water objects (straight-line distance) | ||||
is Distance to the seacoast (straight-line distance) | ||||
is Distance to the town hall (straight-line distance) | ||||
is Aggregating factor: , , (PCA) | ||||
is Distance to the closest hospital (straight-line distance) | ||||
is Distance to the closest nursery school (straight-line distance) | ||||
is Distance to the closest school (straight-line distance) | ||||
Durbin-Watson: | 1.572 | |||
Jarque-Bera (JB): | 3.713 | |||
-value (JB): | 0.156 | |||
Cond. No.: | 2.87 |
@@
We start the modeling by trying the simple and interpretable model – the linear regression. As we mentioned above for the selection of features we apply Forward selection using a correlation matrix (Figure 9). The data was split into two parts – a training set (70% of the data) and a testing set (30% of the data). Table 3 shows the result of the training of the model (for building linear regression we used the “statsmodels” (Seabold and Perktold, 2010) library).
Factors | VIF |
Distance to the water objects (straight-line distance) | 1.463 |
Distance to the seacoast (straight-line distance) | 1.320 |
Distance to the town hall (straight-line distance) | 1.936 |
@@
Aggregate: Distance to the hospitals (straight-line distance) & Distance to the closest nursery school (straight-line distance) & Distance to the closest school (straight-line distance) | 1.424 |
Using this table we can check the model on basic requirements for linear regression. We see that the coefficients of variables are significant (-value 0.01); in general, the equation of regression is significant (-value(-statistic) 0.01); the coefficients of variables have the right sign according to economic substance; model residuals are normally distributed (-value(JB) 0.05) and do not have the strong autocorrelation (1.5 Durbin-Watson 2.5); adjusted coefficient of determination is significant () and it does not differ sufficiently from .
Also factors included in the model do not have multicollinearity because the value of the variance inflation factor (VIF) is less than 10 for each factor (Shrestha, 2020) (Table 4).
@@
Figure 13 shows the distribution of the model residuals. This distribution is approximately normal which satisfies the assumptions of linear regression. Also, we tested the normality of residuals using the library “SciPy”. This test compares calculated -value with the significance threshold (for this paper we chose ). If calculated -value is more than we can assume that the distribution of model residuals is normal (in our case -value = 0.068) (D’Agostino, 1971; D’Agostino and Pearson, 1973).
@@
Also, we created the scatter plot of values of residuals from their numerical order. From Figure 14 we can see that the data does not contain any dependencies. It means that residuals are homoscedastic and the other assumption of linear regression is satisfied.
@@
On the base of the described results, we can state that our model is effective and interpretable. As we have already said the key factor for the identification of the price of a land parcel is its location. We see that linear regression does relatively well but the price of an object can include latent (local-territorial) characteristics. In order to take this into account we tried to apply a powerful instrument dealing with spatially distributed data – geostatistics. Geostatistical methods allow to model the distribution of objects in space. The central task of geostatistics is to reconstruct a studied phenomenon on the base of the values calculated in a limited number of points (Dem’yanov and Savel’eva, 2010). Mathematically we can contemplate this problem as an interpolation problem. For getting the best spatial estimations kriging methods are used. Estimations obtained by kriging have a minimal variation of error. Thus kriging is the best estimator in a class of linear interpolators (Dem’yanov and Savel’eva, 2010). There are two basic conditions for using these methods – data should have spatial continuity and stationarity (Kovalevskiy, 2012). The first condition means that shorter distances between the points should in general lead to closer values of the variable in these points. The second condition means that in some locality the studied phenomenon must be spatially homogeneous.
@@
We can check the condition of stationarity using the probability density function of a variable, in our case we analyze the distribution of residuals of our linear model. If the histogram is similar to the Gaussian distribution then we can assume the stationarity (statistical homogeneity). Strictly speaking, this criteria is not a sufficient condition for confirmation of stationarity but it is strong enough and sufficient condition for practical work (Kovalevskiy, 2012). Figure 13 and the result of the Jarque–Bera test confirm the stationarity in the linear model residuals. It is worth noting that before checking the data on stationarity we need to examine whether the data contains a trend. In our case, we removed the trend from the data using linear regression and then we modeled its residuals with geostatistical methods. The trend in the data can distort the variogram – an instrument for estimation of spatial correlations in data (Dem’yanov and Savel’eva, 2010; Isaaks, 1989). The variogram model is a function describing a dependence between the studied spatial variable and the distance. If data contains a trend then a variogram would reproduce a large-scale trend that leads to missing a correlation of the observed variable on a small scale.
@@
The second condition for implementing kriging – the spatial continuity. We checked it by plotting the correlation of points as a function of the distance between them. From Figure 16 we see that with increasing distance between points, the correlation decreases. Therefore this condition is satisfied.
Thus the final value of the per square meter price is the sum of the trend that we got using the linear regression and the prediction of the linear model residuals that we got using kriging.
@@
OLS Regression Results | ||||
Dep. variable: | ||||
: | 0.319 | |||
: | 0.318 | |||
-statistic: | 227.8 | |||
-value (-statistic): | 1.22-317 | |||
coef | std err | t | ||
1.538+05 | 378.085 | 406.867 | ||
-8730.2980 | 384.482 | -22.707 | ||
3472.7788 | 552.203 | 6.289 | ||
-1075.3009 | 465.506 | -2.310 | ||
-3200.5609 | 445.592 | -7.183 | ||
3374.1161 | 495.161 | 6.814 | ||
4874.6153 | 439.653 | 11.087 | ||
-4124.0290 | 455.890 | -9.046 | ||
-5892.8762 | 384.745 | -15.316 | ||
is Per square meter price | ||||
is Area of flat | ||||
is Distance to the water objects (straight-line distance) | ||||
is Distance to the seacoast (straight-line distance) | ||||
is Distance to the closest railway station (straight-line distance) | ||||
is Distance to the area of negatively influencing factors (straight-line distance) | ||||
is Number of cafes (within a radius of 1 km) | ||||
is Distance to the town hall (distance by the roads) | ||||
is Wall material: large-panel | ||||
Durbin-Watson: | 1.952 | |||
Jarque-Bera (JB): | 791.269 | |||
-value (JB): | 1.51-172 | |||
Cond. No.: | 2.96 |